For the experiment I started with 3.79g of benzyltriphenyl phosphonium chloride and ended up with 1.624g. What is my theoretical and percent yield?

Molecular weight of benzyltriphenyl phosphonium chloride: 388.875 g/mol. Density: 1.18 g/cm^3

No chemical reaction given. Please tell me exactly what you need to figure out the answer. Experiment was "Preparation of (Z)- and (E)-stilbenes by a Wittig Reaction"
Answer #1

Weight of benzyltriphenyl phosphonium chloride = 3.79 g

Number of moles of benzyltriphenyl phosphonium chloride = weight ÷molecular weight

3.79 ÷ 388.875 = 0.00974 moles.

0.00974 moles of benzyltriphenyl phosphonium chloride will give 0.00974 moles of stilbenes

Molecular weight of stilbenes = 180.25 g

The theoretical yield = 0.00974 × 180.25 g . = 1.755 g

Theoretical yield = 1.755 g.

The Percent yield = experimental yield ÷ theoretical yield × 100.

= 1.624÷1.755×100 = 92.53%

Percent yield = 92.53%.